GIVEN THAT P(A)=0.3 AND P(B)=0.5, AND P(A AND B)=0.12, DETERMINE P(A|B)
Answer by robertb(5830)

P(A|B) = P(AB)/P(B) = 0.12/0.5 = 0.24.